OpenStudy (anonymous):

Which set of parametric equations represents the following line or conic? Line: passes through (1,9) and (-6,-1) Use x=x1+t(x2-x1) and y=y1+t(y2-y1)

OpenStudy (anonymous):

7y=10x+53

OpenStudy (anonymous):

x = - 7t + 1 y = - 10t + 9 B. x = - 7t + 1 y = 10t + 9 C. x = - 7t + - 1 y = - 10t - 9 D. x = - 10t + 9 y = - 7t + 1 E. x = - 10t - 9 y = - 7t - 1 these are my answer choices...

OpenStudy (anonymous):

A